# Explaining word embeddings with perfect fidelity: Case study in research impact prediction

### v1

@@ Introduction
Random-split testing. Random, time-sensitive splits of data are a common evaluation technique. In our setup, 80% of data of selected periods is used to train the models and the remaining 20% data from the same periods is used for testing.
Time-split testing. This split is stricter than random-split testing, as the testing set consists of articles that were both published and cited later than the training set in order to prevent any leakage from the test data to the training data.
For training, we use 80% of available articles from publication years $t-1,t-2,\dots$ with the target being derived from citation counts of these articles accumulated by year $t-1$. The test articles were those published in year $t$ with the target being derived from citation counts also from year $t$. This is a particularly challenging setup because the prediction of citations for articles published in the same year is often very difficult.
Semi-split testing. Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest articles, which negatively affects their ability to evaluate the impact of recent research. This semi-split evaluation technique best corresponds to how we envisage that the models would be used by leveraging the latest available citation counts for testing.
Focus on groundbreaking research
To evaluate the ability of the model to separate groundbreaking research from uninfluential articles, we construct a special testing set Extreme10%. Using citation count as a proxy for the impactfulness of the research. This is a subset of the respective test set containing a selection of 5% of articles with the highest citation counts and 5% of articles with the lowest citation counts.
Evaluation measures for the prediction task
Evaluation scenarios
The last column of Table 3 identifies the given combination of publication and citation year used in the train and test data and type of split as the evaluation scenario. For example, for training the classification model in scenario 1 (publication years 2016 and 2017), 80% of all articles published in 2016+2017 with citations in 2017 were used, which makes a total of (962+375+629+368) = 2334 x 0.8 = 1867 articles, which also corresponds to the value in Table 1.
Evaluation of explanations through correlation of feature importance with actual impact
A key element of our research is being able to identify which words or entities (drugs, vaccines, etc.) are linked to high research impact. For a well-performing measure of feature importance, we should be able to observe a positive correlation between feature importance and the presence of the word among articles with observed research impact. High positive feature importance (for the highly cited class) will correlate with the presence of the word among highly cited papers, and vice versa for negative feature importance.
@@
where ##\## denotes the number of elements of a set and $H_{2021}$ is the set of highly-cited articles in 2021 computed as follows:
@@
Given a trained model, we will use the top 100 words with the highest SCOREs (for LR-W2v) or ORCs (for LR-BoW) and bottom 100 words with the lowest values and compare these with the actual impact according to (13). The baseline for this comparison is the probability of an article being highly cited, computed as:
@@
Evaluation of explanations through feature perturbations
To compare the two feature importance methods used in our work, we used a standard approach based on Area Over the Perturbation Curve (AOPC) evaluation [20]. Specifically, we employed AOPCglobal [21], which is a modification of AOPC for benchmarking global feature importance measures. AOPCglobal, unlike the original AOPC, which assesses importance based on local explanations, operates by sequentially removing words based on their global importance as determined by the global feature importance. The SMER method readily provides global importance scores. To obtain these for LIME, which is a local method, we utilized GALE (Global Aggregations of Local Explanations) [21]. We used two variations of GALE: Global LIME Importance and Global Average LIME Importance [21]. Next, features were removed in the order of global importance. The removals were cumulative, first feature, first two features, etc. After each removal, the model was applied to the test set (with removed words), and the average change (typically decrease) in the predicted class probabilities was recorded. Based on this, the AOPCglobal curve was constructed by plotting the number of removed words (X-axis) against the decrease in the class probabilities (Y-axis).
To quantify the aggregate loss in predictive accuracy consequent to the systematic removal of the most significant features, we involved a trapezoidal rule integration, which was used to compute the area under the AOPCglobal curves. The curves were divided into equidistant segments, and areas of individual trapezoids were summed. For the evaluation of global metrics, the curve was segmented into 10 parts, corresponding to each curve consisting of a total of 11 data points (equating to the removal of 0 to 10 words).
Identification of the newly emerged impactful entities
$\textit{CScore}(w)=\textit{SCORE}(w)\times\textit{NumArt}(w),$ (15)
where NumArt(w) is the number of articles where the word appears:
Evaluation of predictions of impactful articles through bibliometric indicators
Bibliometric indicators are widely accepted to correlate with research quality. For example, combined impact factor and citation counts were previously used to predict research impact [22]. As a supplementary approach for validating predictions, we use bibliometric information related to the journal where the article appeared, as well as a bibliometric profile of its authors. We emphasize that this information has not been used in our training setup and is used only for validation.
The quality of the journal is measured through the ISI impact factor, which is computed as the citation counts to articles the journal published in 2019-2020 from articles published in the year 2021 in other journals with an impact factor divided by the number of articles published by the journal in the years 2019-2020. As a reference for a high-impact prediction, we use the median impact factor among the top 10% journals in the category “Health sciences” category (OECD category FORD 3030, data for the year 2021). To represent the research reputation of the authors, we use the maximum of the first and last authors’ Hirsch-index (H-index) values. The H-index is defined as the number of the number of articles published by the author that have at least as many citations as this number. As a reference for a high-impact prediction, we use the median H-index of full professors, which was determined by analyzing the medical education journal editorial boards [23]. Previous research has found that the H-index was a good indicator of citation in the first few years after publication [22].
@@ Results
Experiment 1: Predictions of impactful articles
The goal of this experiment is to evaluate the predictive performance on the task of identifying the most (and least) promising research. We compare the proposed method, logistic regression over word embeddings (further referred to as Word2Vec or w2v), with an older approach based on bag of words (BoW).
To train the models, we created four datasets, each covering articles published in two consecutive years from the period 2016-2020. The overview of the datasets is in Table 3.
For models using embeddings, for each of these datasets, embeddings were trained separately on abstracts published in the corresponding years.
The results are presented in Table 3. Although in terms of random-split evaluation, the BOW method most often has a higher AUC , LR-W2v model outperforms LR-BoW in most of the time-split & semi-split scenarios ($\{3\}$,$\{6\}$,$\{12\}$). The probable reason why in scenario $\bf\{9\}$ both models are failing (as AUC is close to 0.5) is that the training occurred on pre-pandemic data, and articles published during the pandemic were used for testing. The interest of researchers (and consequently the citations) shifted considerably as a result of the pandemic.
On the reference evaluation setup $\{12\}$, which was evaluated based on the most recent set of citations, the best AUC of 0.77 was obtained by the embeddings-based model (trained with word2vec). Overall, the embedding model seems to be better able to adapt to concept drift as measured in the time-split scenarios, where pre-pandemic data were used for training and evaluated on articles that originated subsequently.
Experiment 2: Correlation of feature importance with actual impact
The logistic regression was trained on articles published in 2019-2020 for which citation data were available. The overview of the dataset is in Table 3 (scenario #12). The citation counts from 2021 were used for the validation.The models were trained with and without regularization. We chose the regularized versions (ridge regression) since they performed better in terms of correlation with ActImpact for both models.
Figure 2 answers the question of whether the top 100 words with the highest predicted values, denoted as $T_{100}^{W2V}$ and $T_{100}^{BOW}$, frequently appear in articles that later turned out to be highly cited and whether words with the lowest predicted values, denoted as $L_{100}^{BOW}$ and $L_{100}^{W2V}$, appear in abstracts of articles that later turned out to be lowly cited.
Figure 2: Evaluation of feature importance values for SMER and a baseline method using the Pearson correlation coefficient between the Actual impact of the word and its predicted feature importance. Left: The horizontal axis shows the score of a regularized Logistic Regression model trained on Word2vec embeddings (LR-W2V reg). Right: Reference method based on regularized Logistic Regression Odds Ratio Coefficients for bag of words (LR-BoW reg). Vertical axis indicates the proportion of the articles containing a given word which belongs to the highly cited category. Dots correspond to words $w\in T_{100}^{W2V}\cup L_{100}^{W2V}$ (left) and $w\in T_{100}^{BOW}\cup L_{100}^{BOW}$ (right figure). The red line corresponds to a random baseline ($p_{high}\approx 0.57$).
@@
Experiment 3: Comparing SMER vs LIME through feature perturbations
@@
The previous two experiments showed that word embeddings are superior to the bag-of-words representation both in terms of predictive performance and, when coupled with the proposed SMER method, also when it comes to explanations. The method that we used in Experiment 2 to measure the quality of explanations was based on correlation with actual impact. In this next experiment, we compare SMER with LIME, a frequently used baseline explanation approach.
@@
The data utilized for training the embeddings and the logistic regression model were the same as detailed in Experiment 2, which comprised articles published in years 2019 and 2020, with citation counts from 2021. For the application of both LIME and SMER explanations, we used the logistic regression model trained in Experiment 2.
@@
LIME explainer was configured to use 15 features (with 10 being the default). As a result, the top 15 most influential words were selected for generating explanations for each abstract. We also tried to perform minor adjustments in kernel width or the number of samples, but since this did not significantly alter the outcomes, other parameters were set to their default values.
@@
Figure 3: AOPCglobal for different feature importance metrics. The higher the decrease in predictive performance when top-ranked words are removed (from left to right), the better the feature importance method is in identifying salient features.
@@
Figure 3 shows a summary plot of the AOPCglobal comparison results for LIME, SMER, and a baseline that removed randomly selected words. Table 4 shows the area under the AOPC curve. The random selection, shown by a dashed line, remains almost constant at zero, showing that random word removal has no significant effect on the model’s predictions as expected. An increasing number of removed words results in decreasing changes in the average predicted class probability for both SMER and LIME, which indicates that words associated with the highest scores according to these methods, which were removed first, had the biggest influence. Out of the two aggregation methods for LIME, Global Average LIME Importance had better results; nevertheless, the best overall results were obtained by SMER.
@@
SMER SCORE 0.71
Global Average LIME Importance 0.69
Global LIME Importance 0.33
@@
The previous three experiments validated SMER through the quantitative evaluation of aggregate results using standard methodologies: train-test splits (experiment 1) and feature importance metrics (experiments 2 and 3). The goal of this experiment is to validate SMER by evaluating the utility of specific predictions of significant recently introduced virus strains, drugs and vaccines.
To train the model on the most recent data, we used articles published between 2019 and 2020 as the set of training instances for the logistic regression model. The citation counts were used from the year 2021. Word vectors (embeddings) were trained using the embeddings-based model on recent articles published in the years 2019-2022. The rationale for including the year 2022 in the training of the embeddings was to simulate the situation where the text of articles is available but not their actual impact (citation counts).
@@
(a) words on positions 1-8 by score
Word topic SCORE NumArt CScore
omicron virus 0.998 1,253 1250
b1351 virus 0.972 981 954
b16172 virus 0.998 714 713
b11529 virus 0.999 349 349
vitt condition 0.844 205 173
b1617 virus 0.998 122 122
casirivimab drug 0.996 92 92
covaxin drug 0.958 95 91
(b) words on position 9-15 by score
Word topic SCORE Count CScore
b16171 virus 0.992 91 90
imdevimab drug 0.997 86 86
pango virus 0.994 84 83
covishield vaccine 0.615 130 80
t478k mutation 0.988 73 72
b1526 virus 1.000 69 69
e484q mutation 0.991 69 68
@@
Table 5: Top 15 words sorted by $CScore=SCORE*NumArt$, where SCORE is the predicted impact using SMER (Eq. 4) and NumArt (Eq. 16) is the count of articles containing the word. Topic is a manually assigned semantic category. Words in bold were selected for a follow-up analysis in Table 6.
From Table 5, representatives of each of the three topics were selected by the highest score value and used as seeds to retrieve additional highly-scored words in each of these three domains. This was done by choosing all words whose word-vector similarity (cosine similarity) was at least 70% to the seed word, sorting the words by SCORE, and taking the first 15 hits and excluding the seed word. The results are shown in Table 6(a) (virus variants — words similar to “b11529” - common SARS-CoV-2 variant), Table 6(b) (drugs — similar to “imdevimab”) and Table 6(c) (vaccines — similar to “covaxin”). In addition to the three tables generated automatically by selecting the highest ranked words in each category, we generated a fourth table (Table 6(d)) containing the highest predicted words by similarity to “Remdesivir”. This choice of the seed word followed a previous analysis [24] that detected topics related to COVID-19 and identified the antiviral drug Remdesivir as the dominant term for the drug topic.
An interesting observation is that Table 6(b) contains several terms related to the Moderna vaccine (’moderna’, ’spikevax’, ’modernas’, ’mRNA-1273’). The two main terms from this list (by frequency) are ”moderna” and ”mRNA-1273”, both with a count of about 800 and very similar importance (SCORE) of 0.9548 and 0.9681, respectively. The high agreement in the score for these close synonyms, which, however, have distinct word vectors and were thus handled as two completely separate words, show that the prediction results are consistent. This also indicates that when the count of articles where the term appears is low, the score estimates are less stable, as can be seen with ”spikevax” (has a count of 41) and ”modernas” (count of only 16), which both have lower scores.

### v2

@@ Results
Although in terms of random-split evaluation, the BoW method most often has a higher AUC , LR-W2v model outperforms LR-BoW in most of the time-split & semi-split scenarios ($\{3\}$,$\{6\}$,$\{12\}$). The probable reason why in scenario $\bf\{9\}$, both models are failing as AUC is close to 0.5, is that the training occurred on pre-pandemic data and papers published during the pandemic were used for testing. The interest of researchers (and consequently the citations) shifted considerably as a result of the pandemic.
On the reference evaluation setup $\{12\}$, which was evaluated based on the most recent set of citations, the best AUC of 0.77 was obtained by the embedding-based model (trained with Word2vec). Overall, the embedding model seems to be better able to adapt to concept drift as measured in the time-split scenarios, where pre-pandemic data were used for training and evaluated on papers that originated subsequently.
Experiment 2: Correlation of feature importance with actual impact
The models were trained with and without regularization. We chose the regularized versions (ridge regression) since they performed better in terms of correlation with ActImpact for both models.
Datasets
The logistic regression was trained on papers published in 2019-2020 for which citation data were available. This experiment reuses the same corpus for scenario 2019–2020 already shown in Table 2.
The overview of the dataset is in Table 3 (scenario #12). The citation counts from 2021 were used for the validation.
Evaluation strategy
The explanations are evaluated through the correlation of feature importance with actual impact.
A key element of our research is being able to identify which words or entities (drugs, vaccines, etc.) are linked to high research impact. For a well-performing measure of feature importance, we should be able to observe a positive correlation between feature importance and the presence of the word among papers with observed research impact. High positive feature importance (for the highly cited class) will correlate with the presence of the word among highly cited papers, and vice versa for negative feature importance.
@@
where ##\## denotes the number of elements of a set and $H_{2021}$ is the set of highly-cited papers in 2021 computed as follows:
@@
Given a trained model, we will use the top 100 words with the highest SCOREs (for LR-W2v) or ORCs (for LR-BoW) and bottom 100 words with the lowest values and compare these with the actual impact according to (13). The baseline for this comparison is the probability of a paper being highly cited, computed as:
@@
Figure 2 answers the question of whether the top 100 words with the highest predicted values, denoted as $T_{100}^{W2V}$ and $T_{100}^{BOW}$, frequently appear in papers that later turned out to be highly cited and whether words with the lowest predicted values, denoted as $L_{100}^{BOW}$ and $L_{100}^{W2V}$, appear in abstracts of papers that later turned out to be lowly cited.
Figure 2: This figure shows the evaluation of feature importance values for SMER and a baseline method using the Pearson correlation coefficient between the Actual impact of the word and its predicted feature importance. Left: The horizontal axis shows the score of a regularized Logistic Regression model trained on Word2vec embeddings (LR-W2V reg). Right: Reference method based on regularized Logistic Regression Odds Ratio Coefficients for BoW (LR-BoW reg). Vertical axis indicates the proportion of the papers containing a given word which belongs to the highly cited category. Dots correspond to words $w\in T_{100}^{W2V}\cup L_{100}^{W2V}$ (left) and $w\in T_{100}^{BOW}\cup L_{100}^{BOW}$ (right figure). The red line corresponds to a random baseline ($p_{high}\approx 0.57$).
@@
Experiment 3: Comparing SMER through feature perturbations
@@
The previous two experiments showed that word embeddings are superior to the bag-of-words representation both in terms of predictive performance and, when coupled with the proposed SMER method, also when it comes to explanations. The method that we used in Experiment 2 to measure the quality of explanations was based on correlation with actual impact. In this next experiment, we compare SMER with LIME, SHAP and Global Surrogate Importance methods.
Datasets
@@
The data utilized for training the embeddings and the logistic regression model were the same as detailed in Experiment 2, which comprised papers published in years 2019 and 2020, with citation counts from 2021. For the application of both LIME and SMER explanations, we used the logistic regression model trained in Experiment 2.
Evaluation strategy
To compare the two feature importance methods used in our work, we used a standard approach based on Area Over the Perturbation Curve (AOPC) evaluation [20]. Specifically, we employed AOPCglobal [21], which is a modification of AOPC for benchmarking global feature importance measures. AOPCglobal, unlike the original AOPC, which assesses importance based on local explanations, operates by sequentially removing words based on their global importance.  [^1] Although the name refers to ”Area Over the Perturbation Curve”, the metric is commonly computed as the area under this curve—specifically, under the plot of performance degradation as informative features are removed. This interpretation allows for intuitive comparison: higher area means the model relies more heavily on the top-ranked features, indicating that the importance ranking captures genuinely influential words.
The SMER method readily provides global importance scores based on its internal structure, which can be directly used to rank and remove words for AOPCglobal evaluation.
To obtain these for LIME, which is a local method, we utilized GALE (Global Aggregations of Local Explanations) [21]. We used two variations of GALE: Global LIME Importance and Global Average LIME Importance [21]. SHAP Importance was obtained by computing local SHAP values using the SHAP TextExplainer [22] and then aggregating the absolute contributions of each word across the test set, following the approach for estimating global importance from local explanations.
In addition, we included Global Surrogate Importance, obtained from a decision tree. The probabilistic output of the base model (a value in the 0–1 interval representing the probability of “high impact”) is treated as a continuous target, so we train a regression decision tree on it using TF–IDF (ngrams of length 1–2, vocabulary limited to 5000 terms) inputs to approximate the probability outputs of the embedding-based classifier The surrogate model was trained with a maximum depth of 100 and achieved a high fidelity of $R^{2}=0.89$ on the test set. This approach follows the model extraction framework of Bastani et al. [23], which builds interpretable approximations of complex classifiers for post-hoc explanation purposes.
Next, features were removed in the order of global importance. The removals were cumulative, first feature, first two features, etc. After each removal, the model was applied to the test set (with removed words), and the average change (typically a decrease) in the predicted class probabilities was recorded. Based on this, the AOPCglobal curve was constructed by plotting the number of removed words (X-axis) against the decrease in the class probabilities (Y-axis).
To quantify the aggregate loss in predictive accuracy consequent to the systematic removal of the most significant features, we employed a trapezoidal rule integration, which was used to compute the area under the AOPCglobal curves.
The curves were divided into equidistant segments, and areas of individual trapezoids were summed. For the evaluation of global metrics, the curve was segmented into 10 parts, corresponding to each curve consisting of a total of 11 data points (equivalent to the removal of 0 to 10 words).
Results
@@
LIME explainer was configured to use 15 features (with 10 being the default). As a result, the top 15 most influential words were selected for generating explanations for each abstract.
We also tried to perform minor adjustments in kernel width or the number of samples, but since this did not significantly alter the outcomes, other parameters were set to their default values.
@@
Figure 3: This figure presents the AOPCglobal for different feature importance metrics. The higher the decrease in predictive performance when top-ranked words are removed (from left to right), the better the feature importance method is in identifying salient features.
Figure 3 shows a summary plot of the AOPCglobal comparison results for LIME, SMER, and a baseline that removed randomly selected words. In addition to these methods, we also evaluated SHAP Importance and Global Surrogate Importance.
SHAP Importance was obtained by aggregating local SHAP values across the test set. For each word, we computed the average of its absolute SHAP contributions, resulting in a global importance score.
Global Surrogate Importance was extracted from the trained decision tree by measuring the total reduction in prediction error (mean squared error) attributed to each feature across all splits. These values were then normalized and ranked.
@@
Table 4 shows the area under the AOPC curve. The random selection, shown by a dashed line, remains almost constant at zero, showing that random word removal has no significant effect on the model’s predictions, as expected. An increasing number of removed words results in decreasing changes in the average predicted class probability for all tested methods, which indicates that words associated with the highest scores according to these methods, which were removed first, had the biggest influence. Out of the two aggregation methods for LIME, Global Average LIME Importance had better results; nevertheless, the best overall results were obtained by SMER. SHAP Importance outperformed both Global LIME and Global Surrogate methods, although it did not reach the level of SMER or Global Average LIME Importance.
@@
SMER SCORE 0.50
Global Average LIME Importance 0.48
SHAP importance 0.33
Global surrogate importance 0.27
Global LIME Importance 0.26
@@
The previous three experiments validated SMER through the quantitative evaluation of aggregate results using standard methodologies: train-test splits (experiment 1) and feature importance metrics (experiments 2 and 3). The goal of this experiment is to validate SMER by evaluating the utility of specific predictions of significant recently introduced virus strains, drugs and vaccines. This analysis and interpretation were conducted with significant input and validation from our domain expert in biosciences (M.P.J.) to ensure the biological relevance.
Datasets
To train the model on the most recent data, we used 6,964 papers published between 2019 and 2020 as the set of training instances for the logistic regression model. The citation counts were used from the year 2021.
Word vectors (embeddings) were trained using the embedding-based model on recent papers published in the years 2019-2022 with a total of 414,174 papers. The rationale for including the year 2022 in the training of the embeddings was to simulate the situation where the text of papers is available but not their actual impact (citation counts).
Evaluation strategy
$\textit{CScore}(w)=\textit{SCORE}(w)\times\textit{NumArt}(w),$ (15)
where NumArt(w) is the number of papers where the word appears:
Results
@@
Rank Word Topic SCORE NumArt CScore
1 omicron virus 0.998 1253 1250
2 b1351 virus 0.972 981 954
3 b16172 virus 0.998 714 713
4 b11529 virus 0.999 349 349
5 vitt condition 0.844 205 173
6 b1617 virus 0.998 122 122
7 casirivimab drug 0.996 92 92
8 covaxin drug 0.958 95 91
9 b16171 virus 0.992 91 90
10 imdevimab drug 0.997 86 86
11 pango virus 0.994 84 83
12 covishield vaccine 0.615 130 80
13 t478k mutation 0.988 73 72
14 b1526 virus 1.000 69 69
15 e484q mutation 0.991 69 68
@@
Table 5: Top 15 words sorted by $CScore=\text{SCORE}\times\text{NumArt}$, where SCORE is the predicted impact using SMER (Eq. 5) and NumArt (Eq. 16) is the count of papers containing the word. Topic is a manually assigned semantic category. Words in bold were selected for a follow-up analysis in Table S1.
From Table 5, representatives of each of the three topics were selected by the highest score value and used as seeds to retrieve additional highly-scored words in each domain. We selected all words whose word-vector similarity (cosine $\geq$ 0.70) to the seed exceeded the threshold, sorted them by SCORE, and kept the top 15 hits (excluding the seed itself).
In addition to the three tables generated automatically by selecting the highest ranked words in each category, we generated several additional tables, which are located in the supplementary material:
  • Table 1(a) — virus variants related to the seed “b11529”
  • Table 1(b) — vaccine-related terms similar to “covaxin”
  • Table 1(c) — antibody and drug terms similar to “imdevimab”
  • Table 1(d) — drug terms similar to “remdesivir”